OpenStudy (anonymous):

A bag contains 45 marbles of various colors. Jon draws 60 times, each time replacing the marble before drawing again. He gets a total of 12 blue marbles. Which is best prediction of how many blue marbles are in the bag? A. 5 B. 9 C. 12 D. 15

OpenStudy (anonymous):

I would say a 60/12=5 idk though just what ithink

OpenStudy (zarkon):

let x= number of blue the probability of blue is x/45 the expected value of blue is \(\large {60\frac{x}{45}=x\frac{4}{3}}\) set equal to 12 \[x\frac{4}{3}=12\] \[x=9\]